# Adapting Multi-modal Large Language Model to Concept Drift From Pre-training Onwards

### v2

@@ 2.3 T-distributed Vision Language Model for the Concept Drift
The architecture of the proposed vision language model contains the image encoder, text encoder and text decoder. With an input image $I_{i}\in\mathbb{R}^{H\times W}$, the visual features $x_{\text{img}}$ are extracted by the image encoder $E_{\text{img}}$ and further projected to the spherical eigenspace by the L2 normalizer $P_{\text{norm}}$:
$x_{\text{img}}=P_{\text{norm}}(E_{\text{img}}(I_{i}))\in\mathbb{R}^{n\times d}$ (7)
@@
where $n$ is the number of visual features and $d$ represents the feature dimension. The image encoder $E_{\text{img}}$ can be any common visual backbones, such as Vit-Base [9], Vit-Large [9] and ResNeXt-50 [10]. In terms of the text encoder $E_{\text{txt}}$, with a processed input text sequence $T_{i}$, text features are extracted by the language encoder and further projected to the spherical eigenspace by the L2 normalizer $P_{\text{norm}}$:
$x_{\text{txt}}=P_{\text{norm}}(E_{\text{txt}}(T_{i}))\in\mathbb{R}^{n\times d}$ (8)
where $n$ is the number of input tokens and $d$ represents the feature dimension. In our case, Bert [11] is used as the language encoder.
@@
In the pre-training of the vision language model, the T-distributed adapter aligns image and text features by contrastive learning. It seeks to align visual and textual transformer feature spaces by promoting similar representations for positive pairs and dissimilar representations for negative pairs. More importantly, our approach circumvents model bias stemming from the long-tailed distribution of data. It couples with the language modelling loss to drives the training of all three modules.
@@
In the downstream classification task, we leverage the T-distributed adapter as the router to distribute features to various FFNs as experts. Furthermore, by explicitly modelling the feature space, the T-router enables a straightforward application of non-parametric KNN to effectively partition the data into ID and OOD samples. Besides, following the Blip [12], the image encoder and the text decoder are frozen out of training during fine-tuning. A head of the classifier with two linear layers embeds features into the spherical eigenspace is trained.
@@ 2.4 Building Multi-modal Dataset OpenMMlo for the Long-Tailed Open World
Due to the inherent challenge of obtaining images and related captions, most multi-modal datasets struggle to follow the long-tailed distribution in the open world. Recognizing the demand for higher-quality multi-modal data with long-tailed distribution in an open world, we developed a group of datasets called Open Multi-modal Long-Tailed OOD Datasets (OpenMMlo).
We extend the open-source datasets, namely ImageNet-LT [13], iNatualist2018 [14] and Places-LT [13]. ImageNet-LT has 1,000 classes and contains 115.8k samples, with a maximum of 1,280 samples and a minimum of 5 samples for a category. Besides, it consists of 18k images for OOD detection. Places-LT has 184.5K samples from 365 classes, with class samples ranging from 4,980 to 5. The iNaturalist 2018 is a large-scale species dataset collected in the natural world with 437.5K samples for 8,142 classes. We use the InstructBLIP [15] to generate the related caption of the image, with the prompt of "What does this picture describe? Please describe in detail its size, location, color, and its relationship to the surroundings.". For more details about OpenMMlo, please refer to Appendix LABEL:appendix:details.
@@ 3 Experiments
In this section, we first present the performance in downstream long-tailed classification and OOD detection tasks, which is induced by tail drift and OOD drift on MLLMs. Then, we evaluate the interior feature space of the VL model and further demonstrate our method alleviates the crowding and bias problems caused by the tail drift and OOD drift. The constructed long-tailed multi-modal datasets OpenMMlo is utilized for training and validating. In terms of the OOD drift detection, we follow the setting of CIDER [16]. The model is trained on CIFAR100-LT [17] with an imbalance ratio of 100, and validated on external OOD datasets including SVHN [18], Places365 [19], LSUN [20], iSUN [21] and Texture [22]. More detailed experimental implementations are given in Appendix LABEL:appendix:expdetails.
@@ 3.1 Taming the Tailed Drift and OOD Drift for Robust Fine-tuning
We compare our proposed vision-language model with other models to explicitly demonstrate its superior performance in long-tailed open-world scenarios. As shown in Table 1, our model demonstrates exceptional overall performance in long-tailed classification across two large-scale datasets, namely the ImageNet-LT and iNaturalist 2018. Especially, training from scratch means that our method uses the imbalanced dataset for pre-training instead of utilizing the pre-trained model, such as clip [23] pre-trained by the large WIT dataset. The results validate the robustness of our vision language model against biases arising from tailed drift, particularly when leveraging large-scale data for both pre-training and fine-tuning.
Table 1: Evaluation results of long-tailed classification on ImageNet-LT and iNatualist2018. The best-performing models are highlighted in red. Many, Medi. and Few denote the evaluated splits of many-shot ($>$100 training samples), medium-shot (20-100 samples) and few-shot ($<$20 samples). Top-1 accuracy is applied to evaluate the performance of different methods.
@@
cRT [24] 61.8 46.2 27.3 49.6 69.0 66.0 63.2 65.2
LWS [24] 60.2 47.2 30.3 49.9 65.0 66.3 65.5 65.9
MiSLAS [25] 62.9 50.7 34.3 52.7 73.2 72.4 70.4 71.6
BALMS [26] 64.1 48.2 33.4 52.3 - - - 70.6
LADE [27] 64.4 47.7 34.3 52.3 - - - 69.3
ACE [28] - - - 56.6 - - - 72.9
RIDE [29] 68.2 53.8 36.0 56.9 70.9 72.4 73.1 72.6
PaCo [30] 68.2 58.7 41.0 60.0 70.3 73.2 73.6 73.2
NCL [31] ResNet-50 - - - 57.4 72.0 74.9 73.8 74.2
ViT [9] 73.7 46.5 15.6 52.4 65.4 55.3 50.9 54.6
MAE [32] 74.7 48.2 19.4 54.5 79.6 70.8 65.0 69.4
DeiT [33] 70.4 40.9 12.8 48.4 72.9 62.8 55.8 61.0
LiVT [34] ViT-B/16 73.6 56.4 41.0 60.9 78.9 76.5 74.8 76.1
Ours 77.2 68.6 51.3 69.6 83.3 82.1 80.5 81.5
@@
LPT [35] - - - - 62.1 76.2 79.3 76.1
BALLAD [36] 79.1 74.5 69.8 75.7 - - - -
Decoder [37] - - - 73.2 - - - 59.2
VL-LTR [38] 84.5 74.6 59.3 77.2 - - - 81.0
LIFT [39] ViT-B/16 (w/ Pretrained Clip) 80.2 76.1 71.5 77.0 72.4 79.0 81.1 79.1
Ours 79.9 76.8 74.5 77.6 84.1 82.7 81.0 82.1
Compared to other methods trained from scratch, especially the ViT model, our model demonstrates a notable lead across all metrics, indicating our effective mitigation of concept drift during the pre-training, and providing robust pre-trained models for downstream tasks. Furthermore, to compare the current long-tailed methods, we the same setup as LIFT[39], i.e., using the pre-trained model of clip and only fine-tuning. The superior results on medium and few splits of ImageNet-LT demonstrate the adaptability and robustness of our model in dealing with the gradual drift caused by tail data. Besides, although we are slightly behind LIFT[39] in the few split of iNatualist2018, we still surpass it overall, exhibiting that our method does not compromise the accuracy of the head category to improve the tailed.
@@
ProxyAnchor [40] 87.2 82.4 37.2 91.7 70.0 85.0 65.6 85.0
CE + SimCLR [41] 24.8 94.5 56.4 89.0 66.5 83.8 63.7 82.0
CSI [42] 44.5 92.7 75.6 83.8 76.6 85.0 61.6 86.5
SSD+ [43] 31.2 94.2 79.4 85.2 80.9 84.1 66.6 86.2
KNN+ [8] 39.2 92.8 49.0 89.3 75.0 82.7 57.2 88.4
CIDER [16] 12.6 97.8 30.2 92.8 46.0 88.9 35.6 92.3
@@ 3.2 Concept Drift-Aware Image-Text Alignment for Effective Pre-training
Moreover, we verified at the feature-level that the proposed T-distributed adapter significantly alleviates the bias from tailed drift and OOD drift in the pre-training. As exhibited in Table 3, the degree of ID intra-class compactness reduces from 49.2 in LT/cosine to 36.2 in LT/Thp. It thereby validates the effectiveness of the proposed T-distributed adapter in enhancing feature extraction in long-tailed scenarios. Notably, the decrease in standard deviation demonstrates that the model considerably mitigates the bias induced by tail drift. In addition, our method achieves remarkable inter-class separability within in-distribution categories under long-tailed scenarios, even surpassing the performance of cosine achieved under the balanced dataset. It confirms the effectiveness of the proposed high kurtosis method in enhancing the alignment between images and text in the pre-training stage. Concerning the separability between ID and OOD, we achieve superior results even than the balanced condition, attributed to the inherent light-tailed property of the T-distributed adapter. It ensures our approach performs robust for OOD drift detection based on the spherical distribution robustly.
@@ 3.3.1 T-distributed Spherical Embedding in the Pre-training and Fine-tuning
69.6
3.3.2 Various Concentration Parameter $\kappa$
Furthermore, we conduct ablation experiments to examine the impact of different concentrations of the T-distributed adapter on the overall performance of the vision language model in Table 5. Four fixed degrees of concentration are involved, namely $\kappa=4$ 16, 64 and 128. The greater the degree of concentration, the greater the kurtosis of the T-distributed spherical metric. Besides, the concentration can also be utilized as a trainable parameter joining in the training, with the initial setting of 16.
Table 5: Ablation evaluation results of various concentration parameter $\kappa$ on long-tailed classification task. "Training" denotes the $\kappa$ involved in the training as a parameter of the model with an initial setting of 16. The results are based on the ImageNet-LT with the Vit-base. Top-1 accuracy (Acc) is used as the metric.

### v3

@@ 2.3 T-distributed Vision Language Model for the Concept Drift
The proposed vision language model follows the encoder-decoder mixture architecture of the Blip Li et al. (2022c), containing an image encoder, a text encoder and an image-grounded text decoder. With an input image $I_{i}\in\mathbb{R}^{H\times W}$, the visual features $x_{\text{img}}$ are extracted by the image encoder $E_{\text{img}}$ and further projected to the spherical eigenspace by the $L_{2}$ normalizer $P_{\text{norm}}$:
$x_{\text{img}}=P_{\text{norm}}(E_{\text{img}}(I_{i}))\in\mathbb{R}^{n\times d},$ (7)
@@
where $n$ is the number of visual features and $d$ represents the feature dimension. The image encoder $E_{\text{img}}$ can be any common visual backbones, such as Vit-Base Dosovitskiy et al. (2021), Vit-Large Dosovitskiy et al. (2021) and ResNeXt-50 Xie et al. (2017). In terms of the text encoder $E_{\text{txt}}$, with a processed input text sequence $T_{i}$, text features are extracted by the language encoder and further projected to the spherical eigenspace by the $L_{2}$ normalizer $P_{\text{norm}}$:
$x_{\text{txt}}=P_{\text{norm}}(E_{\text{txt}}(T_{i}))\in\mathbb{R}^{n\times d},$ (8)
where $n$ is the number of input tokens and $d$ represents the feature dimension. In our case, Bert Devlin et al. (2019a) is used as the language encoder, where a [CLS] token is added to the start of the text input for sentence summarization. Additionally, an image-grounded text decoder is employed to produce a textual description corresponding to a provided image. Utilizing the input visual features $x_{img}$ and text features $x_{txt}$, we initially create fused multi-modal representations by merging the image and text feature embeddings. These combined features act as the keys and values within the cross-attention blocks in the image-grounded text decoder. Through conditioning on the already predicted partial sequence $y_{i<j}$, the decoder iteratively forecasts the token at position $j$, effectively producing textual descriptions corresponding across modalities.
@@
In the pre-training of the vision language model, the T-distributed adapter aligns image and text encoders by contrastive learning. It seeks to align visual and textual transformer feature spaces by promoting similar representations for positive pairs and dissimilar representations for negative pairs. More importantly, our approach circumvents model bias stemming from the long-tailed distribution of data. Specifically, given a mini-batch with $N$ image-text feature pairs, we calculate the $N\times N$ Thp similarity of the cross between image and text features. $N$ correct pairs are recognized as positive samples to maximize the Thp similarity, whereas the rest of $N^{2}-N$ are negative samples to minimize the similarity. And we follow the ALBEF Li et al. (2021) to use soft labels from a momentum encoder as training targets to account for the potential positives in the negative pairs.
Additionally, coupled with the T-distributed adapter, language modeling loss is utilized to activate the image-grounding text encoder for generating coherent and detailed captions based on the image, further propelling the training of all three modules. Driven by language modeling loss, the model is trained to optimize a cross-entropy loss with label smoothing, to maximize the likelihood of the generated text in an autoregressive manner.
@@
In the downstream classification task, we leverage the T-distributed adapter as the router to distribute features to various FFNs as experts. Furthermore, by explicitly modeling the feature space, the T-router enables a straightforward application of non-parametric KNN to effectively partition the data into ID and OOD samples. Besides, following the Blip Li et al. (2022c), the image encoder and the text decoder are frozen out of training during fine-tuning. A head of the classifier with two linear layers embeds features into the spherical eigenspace is trained.
@@ 2.4 Building Multi-modal Dataset OpenMMlo for the Long-Tailed Open World
As the parameters of large models continue to expand, the demand for extensive training data also escalates. However, due to the inherent challenge of obtaining images and related captions, most multi-modal datasets struggle to be balanced in an open world, while cleaning the data requires huge costs. Thus, our aspiration is for the model to adeptly acclimate to the imbalanced dataset by itself, acquiring abundant knowledge with more and more data but not exhibiting bias. In this context, a more realistic training dataset for vision language models is required to validate their potential to be trained under the long-tailed open world. Recognizing the demand for higher-quality multi-modal data with long-tailed distribution in an open world, we developed a group of datasets called Open Multi-modal Long-Tailed OOD Datasets (OpenMMlo).
We extend the open-source datasets, namely ImageNet-LT Liu et al. (2019), iNatualist2018 Van Horn et al. (2018) and Places-LT Liu et al. (2019). ImageNet-LT has 1,000 classes and contains 115.8k samples, with a maximum of 1,280 samples and a minimum of 5 samples for a category. Besides, it consists of 18k images for OOD detection. Places-LT has 184.5K samples from 365 classes, with class samples ranging from 4,980 to 5. The iNaturalist 2018 is a large-scale species dataset collected in the natural world with 437.5K samples for 8,142 classes. We use the InstructBLIP Dai et al. (2023) to generate the related caption of the image, with the prompt of ”What does this picture describe? Please describe in detail its size, location, color, and its relationship to the surroundings.”. And, we define long-tailed data in image-caption pairs according to the image categories, which are provided in open-source image datasets. Concerning related captions based on images, we counted the word frequencies and found that their distribution is similar to the image categories distribution, which is imbalanced. For more details about OpenMMlo, please refer to Appendix A.4.
@@ 3 Experiments
In this section, we first present the performance in downstream long-tailed classification and OOD detection tasks, which is induced by tail drift and OOD drift on MLLMs. Then, we evaluate the interior feature space of the VL model and further demonstrate our method alleviates the crowding and bias problems caused by the tail drift and OOD drift. The constructed long-tailed multi-modal dataset OpenMMlo is utilized for training and validating. In terms of the OOD drift detection, we follow the setting of CIDER Ming et al. (2023). The model is trained on CIFAR100-LT Krizhevsky & Hinton (2009) with an imbalance ratio of 100, and validated on external OOD datasets including SVHN Netzer et al. (2011), Places365 Zhou et al. (2017), LSUN Yu et al. , iSUN Xu et al. and Texture Cimpoi et al. (2014). More detailed experimental implementations are given in Appendix A.3.
@@ 3.1 Taming the Tailed Drift and OOD Drift for Robust Fine-tuning
We compare our proposed vision-language model with other models to explicitly demonstrate its superior performance in long-tailed open-world scenarios. As shown in Table 1, our model demonstrates exceptional overall performance in long-tailed classification across two large-scale datasets, namely the ImageNet-LT and iNaturalist 2018. To ensure a more equitable comparison, we opted to conduct pre-training using ImageNet and iNaturalist datasets separately, rather than pre-training the entire OpenMMlo. Besides, it is worth noting that training from scratch means that our method uses the imbalanced dataset for pre-training instead of utilizing the pre-trained model, such as clip Radford et al. (2021) pre-trained by the large WIT dataset. The results validate the robustness of our vision language model against biases arising from tailed drift, particularly when leveraging large-scale data for both pre-training and fine-tuning.
As shown in Table 1, compared to other methods trained from scratch, especially the ViT model, our model demonstrates a notable lead across all metrics, indicating our effective mitigation of concept drift during the pre-training, and providing robust pre-trained models for downstream tasks. Furthermore, to compare the current long-tailed methods, we apply the same setup as LIFT Shi et al. (2024), i.e., using the pre-trained model of the clip and only fine-tuning. Only fine-tuning means that the method does not pre-train the model on the long-tail dataset, while directly using the parameters of CLIP pre-trained on high-quality and large-scale WIT dataset. And they only fine-tune the model on long-tailed datasets. It is worth noting that, most vision language models only focus on the impact of tailed drift in the fine-tuning, such as LPT DONG et al. (2023), BALLAD Ma et al. (2021), Decoder Wang et al. (2024), VL-LTR Tian et al. (2022) and LIFT Shi et al. (2024). We are the pioneers in revealing the unexplored impacts of concept drift from pre-training onwards. The superior results on medium and few splits of ImageNet-LT demonstrate the adaptability and robustness of our model in dealing with the gradual drift caused by tail data. Besides, although we are slightly behind LIFT Shi et al. (2024) in the few split of iNatualist2018, we still surpass it overall, exhibiting that our method does not compromise the accuracy of the head category to improve the tailed.
Table 1: Evaluation results of long-tailed classification on ImageNet-LT and iNatualist2018. The best-performing models are highlighted in red. Many, Medium and Few denote the evaluated splits of many-shot ($>$100 training samples), medium-shot (20-100 samples) and few-shot ($<$20 samples). Top-1 accuracy is applied to evaluate the performance of different methods. Additionally, $\dagger$ means the model is trained with the resolution of $384\times 384$. Besides, ZS denotes the zero-shot results of the CLIP model, LP represents the linear probing results, and FT means the fine-tuning results.
@@
cRT Kang et al. (2019) 61.8 46.2 27.3 49.6 69.0 66.0 63.2 65.2
LWS Kang et al. (2019) 60.2 47.2 30.3 49.9 65.0 66.3 65.5 65.9
MiSLAS Zhong et al. (2021) 62.9 50.7 34.3 52.7 73.2 72.4 70.4 71.6
BALMS Ren et al. (2020) 64.1 48.2 33.4 52.3 - - - 70.6
LADE Huang et al. (2016) 64.4 47.7 34.3 52.3 - - - 69.3
ACE Cai et al. (2021) - - - 56.6 - - - 72.9
RIDE Wang et al. (2020) 68.2 53.8 36.0 56.9 70.9 72.4 73.1 72.6
PaCo Cui et al. (2021) 68.2 58.7 41.0 60.0 70.3 73.2 73.6 73.2
NCL Li et al. (2022b) ResNet-50 - - - 57.4 72.0 74.9 73.8 74.2
ViT Dosovitskiy et al. (2021) 50.5 23.5 6.9 31.6 65.4 55.3 50.9 54.6
MAE He et al. (2022) 74.7 48.2 19.4 54.5 79.6 70.8 65.0 69.4
DeiT Touvron et al. (2022) 70.4 40.9 12.8 48.4 72.9 62.8 55.8 61.0
LiVT Xu et al. (2023) 73.6 56.4 41.0 60.9 78.9 76.5 74.8 76.1
LiVT $\dagger$ Xu et al. (2023) ViT-B/16 76.4 59.7 42.7 63.8 83.2 81.5 79.7 81.0
Ours 76.4 66.2 48.9 67.6 82.5 79.8 77.1 78.9
[…]
@@
CLIP+ZS Radford et al. (2021) 82.0 70.4 69.6 70.5 9.9 5.3 4.6 5.5
CLIP+LP Radford et al. (2021) 87.3 65.1 19.0 67.4 62.4 7.1 0.1 10.0
CLIP+FT Radford et al. (2021) 83.0 65.0 39.9 68.5 79.4 67.6 59.1 65.4
LPT DONG et al. (2023) - - - - 62.1 76.2 79.3 76.1
BALLAD Ma et al. (2021) 79.1 74.5 69.8 75.7 - - - -
Decoder Wang et al. (2024) - - - 73.2 - - - 59.2
VL-LTR Tian et al. (2022) 84.5 74.6 59.3 77.2 - - - 81.0
LIFT Shi et al. (2024) ViT-B/16 (w/ Pretrained Clip) 80.2 76.1 71.5 77.0 72.4 79.0 81.1 79.1
Ours 79.5 76.5 74.1 77.2 83.5 82.2 80.7 81.7
Ours $\dagger$ 79.9 76.8 74.5 77.6 84.1 82.7 81.0 82.1
Beyond that, we compare our method with the CLIP Radford et al. (2021) under zero-shot, linear probing and fine-tuning, where CLIP results are from the Decoder Wang et al. (2024). Based on the zero-shot results, it is evident that CLIP, even trained on large-scale and high-quality WIT datasets, struggles to address the issue of tailed drift. CLIP only achieves 5.5% accuracy on iNaturalist2018, and the accuracy variance between many-shot and medium-shot scenarios is 11.6% on ImageNet-LT. Our method significantly outperforms CLIP in dealing with tailed drift, especially on iNaturalist2018. It also indicates that training on a high-quality balanced dataset alone cannot effectively mitigate the bias induced by long-tail drift. Furthermore, the results of linear probing and fine-tuning demonstrate that imbalanced datasets can induce pronounced tailed drift in MLLMs, ultimately degrading model performance. The CLIP accuracy in Few-shot is only 39.9% under fine-tuning on ImageNet-LT, much lower than the 69.6% under zero-shot. It further verifies the challenges brought by imbalanced data in the training of MLLMs and the superiority of our method in adapting the MLLM to concept drift from pre-training onwards.
@@
ProxyAnchor Kim et al. (2020) 87.2 82.4 37.2 91.7 70.0 85.0 65.6 85.0
CE + SimCLR Winkens et al. 24.8 94.5 56.4 89.0 66.5 83.8 63.7 82.0
CSI Tack et al. (2020) 44.5 92.7 75.6 83.8 76.6 85.0 61.6 86.5
SSD+ Sehwag et al. (2021) 31.2 94.2 79.4 85.2 80.9 84.1 66.6 86.2
KNN+ Sun et al. (2022) 39.2 92.8 49.0 89.3 75.0 82.7 57.2 88.4
CIDER Ming et al. (2023) 12.6 97.8 30.2 92.8 46.0 88.9 35.6 92.3
@@
Table 3: Evaluation results of generalization on ImageNet-Sketch Wang et al. (2019) with ImageNet Russakovsky et al. (2015a) as the source dataset. We compare our methods with other VL models, including zero-shot CLIP Radford et al. (2021), linear probing CLIP Radford et al. (2021), CoOp Zhou et al. (2022), VPT Jia et al. (2022) and DAPT Cho et al. (2023).
CLIP+ZS CLIP+LP CoOp VPT DAPT Ours
46.1 36.0 47.1 47.7 48.3 50.2
Moreover, we evaluate the generalizability of our method in the domain generalization setting. Experiments are conducted on ImageNet-Sketch Wang et al. (2019) with ImageNet Russakovsky et al. (2015a) as the source dataset, as shown in the Table 3. From the experiment results, our method achieves superior performance with an accuracy of 50.2% on ImageNet-Sketch, attributed to the robustness of the T-distribution-based drift adapter. It further verifies the generalization ability of our model in the open world.
@@ 3.2 Concept Drift-Aware Image-Text Alignment for Effective Pre-training
Moreover, we verified at the feature level that the proposed T-distributed adapter significantly alleviates the bias from tailed drift and OOD drift in the pre-training. As exhibited in Table 4, the degree of ID intra-class compactness reduces from 49.2 in LT/cosine to 36.2 in LT/Thp. It thereby validates the effectiveness of the proposed T-distributed adapter in enhancing feature extraction in long-tailed scenarios. Notably, the decrease in standard deviation demonstrates that the model considerably mitigates the bias induced by tail drift. In addition, our method achieves remarkable inter-class separability within in-distribution categories under long-tailed scenarios, even surpassing the performance of cosine achieved under the balanced dataset. It confirms the effectiveness of the proposed high kurtosis method in enhancing the alignment between images and text in the pre-training stage. Concerning the separability between ID and OOD, we achieve superior results even than the balanced condition, attributed to the inherent light-tailed property of the T-distributed adapter. It ensures our approach performs robustly for OOD drift detection.
@@ 3.3.1 T-distributed Spherical Embedding in the Pre-training and Fine-tuning
69.4
3.3.2 Various Concentration $\kappa$ in T-Adapter
Table 6: Ablation evaluation results of various concentrations of parameter $\kappa$ on the long-tailed classification task. ”Training” denotes the $\kappa$ involved in the training as a parameter of the model with an initial setting of 16. The results are based on the ImageNet-LT with the Vit-base. Top-1 accuracy (Acc) is used as the metric.